40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Mar 27 04:36:46 2015, den, Update, LSC, als->pdh transition, prcl on 1f, alignment 
    Reply  Fri Mar 27 14:47:57 2015, Koji, Update, LSC, als->pdh transition, prcl on 1f, alignment CARM_transition.pdf
Message ID: 11177     Entry time: Fri Mar 27 04:36:46 2015     Reply to this: 11179
Author: den 
Type: Update 
Category: LSC 
Subject: als->pdh transition, prcl on 1f, alignment 

Tonight I have modified transition steps from als to pdh signals. I have added 1:20 filters to CARM_A and DARM_A filter banks to make them unconditionally stable. These filters made locking more robust -- duty cycle is was ~70% tonight. I have also modified slow/ao crossover to avoid ringing up of lines above 1kHz.

Once AO is engaged with high bandwidth, REFL55 signal looks good and I transition PRCL from 165I to 55I. Optical gain compared to PRMI reduced from 55I/165I = -330 down to 55I/165I = 30 in full lock.

I worked on alignment of ETMs. Looking on the cameras I could improve arm power up to 160 and ifo visibility was 80%. POP22 fluctuated by ~50% and every few minutes we loose lock because POP22 almost touches zero.

ELOG V3.1.3-